OpenStudy (anonymous):

Find the inverse of the function. F(x)=6-3x

OpenStudy (anonymous):

So far I have f(x)=x-6/-3... I need it in y=mx+b.

OpenStudy (anonymous):

@DangerousJesse @Nnesha

OpenStudy (anonymous):

F(x) = y y=x-6/3 3y=6-x 3y-6=-x x=6-3y F'(x) = 6 -3x

OpenStudy (anonymous):

 heres what i have..

OpenStudy (mathmath333):

\(\large y=6-3x\) \(\large y-6=-3x\) \(\large 3x=6-y\) \(\large x=\dfrac{6-y}{3}\) \(\large hence~~f'(x)=\dfrac{6-x}{3}\)

OpenStudy (anonymous):

but how can i get rid of the 3 in the denominator? can i still verify the inverse with the 3, or will i have to do something with it?

OpenStudy (mathmath333):

it can be written as further \(\large f'(x)=2-\dfrac{x}{3}\)

OpenStudy (mathmath333):

nothing can be done with 3 ,it resides

OpenStudy (anonymous):

okay, thanks!

OpenStudy (mathmath333):

welcome